OpenStudy (anonymous):

If three points (x1,y1),(x2,y2) and (x3,y3) lie on the same line,prove that [y2-y3/x2x3 ]+[ y3 -y1 /x3x1]+[y1-y2/x1x2 ]= 0

OpenStudy (stamp):

\[\frac{y_{2}-y_{3}}{x_{2}x_{3}}+\frac{y_{3}-y_{1}}{x_{3}x_{1}}+\frac{y_{1}-y_{2}}{x_{1}x_{2}}=0\]

OpenStudy (anonymous):

how to prove this.